OpenStudy (anonymous):

write in vertex form y=x^2-8x+16-16

OpenStudy (anonymous):

y=a(x-h)^2+k h=-b/2a = 8/2 =4 y=(x-4)^2+k you find k now

OpenStudy (chihiroasleaf):

vertex form is y = a(x-h)^2 + k so, you must modify the expression on the right side hint: (a-b)^2 = a^2 - 2ab + b^2

OpenStudy (anonymous):

would it be 16